OpenStudy (anonymous):

for #1 part b of hw 11, when I find

OpenStudy (anonymous):

for #1 part b of hw 11, when I find \[g _{\omega}=g _{e}g _{7}^{-1}\] but when I go to find \[g _{h}= g _{w}^{-1}g _{e}g _{7}^{-1}\] I always get the 4x4 identity matrix

OpenStudy (anonymous):

\[g _{h}\] is suppose to be a pue rotation matrix. What am I doing wrong?

OpenStudy (anonymous):

How do you find gw?

OpenStudy (anonymous):

\[g _{w} = g _{e}g _{7}^{-1} = g _{i}g _{7}^{-1}\] I am given that l3 = 0.5 so \[g _{7}=\left[\begin{matrix}1 & 0 & 0.5\\ 0 & 1 & 0 \\ 0 & 0 & 1\end{matrix}\right]\] and so \[g _{7}^{-1}=\left[\begin{matrix}1 & 0 & -0.5\\ 0 & 1 & 0 \\ 0 & 0 & 1\end{matrix}\right]\] I am given \[g _{i}\] then I multiply them together to get \[g _{w}\]

OpenStudy (anonymous):

Ah ok you missed an important point... pos(gw) = pos(ge*g7^-1), but gw does not equal ge*g7^-1. You need to get gw by first getting alpha1, alpha2, and alpha3

OpenStudy (anonymous):

I thought Pos(gw) was (xw, yw, zw) and I need alpha 1, 2 and 3 to get xw, yw and zw or visa versa

OpenStudy (anonymous):

Pos(gw) is (xw, yw, zw). Just like you did last week, you use this position to solve for the angles before the wrist. Pos(gw) can be figured out from known quantities.

OpenStudy (anonymous):

I am to use \[\beta\] \[\gamma \] \[\delta \] and r to find \[\alpha _{1}\] and \[\alpha _{2}\]

OpenStudy (anonymous):

I have to go for now be back later. You need to use geometry. It should be in the notes.

OpenStudy (anonymous):

thanks later

OpenStudy (anonymous):

Were you able to move along?

OpenStudy (anonymous):

no, I looked at the geometry section of the elbow manipulator in my notes. All it says is to solve (x,y) projection problem to get alpha1 and the (r,z) projection problem to get alpha2 and alpha3.

OpenStudy (anonymous):

I know what a projection onto a plane means but I do not see how I can get the needed information without having the angle or the x y projections.

OpenStudy (anonymous):

Well you have xw, yw, and zw right?

OpenStudy (anonymous):

I have the \[d _{i}\] which comes from the \[g _{i}\] that is given. I need the \[g _{\omega} \] to get the \[d _{\omega} \] which is \[\left(\begin{matrix}x _{\omega} \\ y _{\omega} \\ z _{\omega}\end{matrix}\right) \] I know that the way I was getting the \[g _{\omega} \] is not right but I do not know of another way.

OpenStudy (anonymous):

How are you getting gw?

OpenStudy (anonymous):

Not to sound short but I thought this would be quicker. I explained it in my forth posting.

OpenStudy (anonymous):

I know now from what you stated that Pos(gw) = Pos(ge*g7^-1) does not mean gw = ge * g7^-1

OpenStudy (anonymous):

Ok let me make it clear. \[pos(g_w) = pos(g_e*g_7^{-1})\]\[g_e = \left[\begin{matrix}R_e & d_e \\ 0 & 1\end{matrix}\right]\]\[g_7 = \left[\begin{matrix}I & \left(\begin{matrix}0 \\ l_3 \\ 0\end{matrix}\right) \\ 0 & 1\end{matrix}\right]\]Therefore\[d_w = \left(\begin{matrix}x_w \\ y_w\\z_w\end{matrix}\right) = pos(g_e*g_7^{-1})\]

OpenStudy (anonymous):

This is exactly the same thing as you did last week but in 3D.

OpenStudy (anonymous):

I see a great similarity in this to what I did. Your saying I do not know ge and that is why you have written ge in its general form. I am thinking that I do know ge because it is gi. I have put l3 in the wrong place. I put it in the x position, I did not see the lettering on the axis in the problem. I do see were l3 should be in the y position.

OpenStudy (anonymous):

I'm not saying ge is unknown I'm just giving you the general solution. In 1b, gi = ge for the first case and gf=ge in the second case.

OpenStudy (anonymous):

In inverse kinematics problem ge is always known. It doesn't make sense to calculate joint angles if you don't know where your end effector is.

OpenStudy (anonymous):

I was looking back at last weeks hw and saw something that looked like what I have in this problem. That I can find gw by \[g _{\omega}=g _{1}(\alpha _{1})g _{2}(\alpha _{2})g _{3}(\alpha _{3})g _{4}\]

OpenStudy (anonymous):

After you get alpha1, alpha2, and alpha3, you can use this to get gw yes.

OpenStudy (anonymous):

That is what I need to finish that problem.

OpenStudy (anonymous):

I did move to the last problem and I have a question.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I am asked to design a splined trajectory from alpha0 to alpha1 and the movement time is to be 8 seconds. I looked in my notes and I see \[A = X ^{-1}\alpha \] but I am not sure how to code it to move in a curved path.

OpenStudy (anonymous):

Do you know how to write a function that would give you the angle as a function of time?

OpenStudy (anonymous):

a(1) = q0; a(2) = qv0; a(3) = (3 * (q1-q0) - t * (2 * qv0 + qv1)) / t^2; a(4) = (-2 * (q1-q0) + t * (qv0 + qv1)) / t^3; q0 is initial position q1 is final position qv0 is initial velocity qv1 is final velocit t is time I took the equations from the notes and solved for alpha

OpenStudy (anonymous):

As for a matlab function, I am not sure.

OpenStudy (anonymous):

Ok I have to go for now. I'll do an example and talk a bit about the matlab in class.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

When I calculate \[g _{h} = g _{\omega}^{-1}g _{e}g _{7}^{-1} = g _{\omega}^{-1}g _{i}g _{7}^{-1} \] I am suppose to get a pure rotational matrix

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I did not

OpenStudy (anonymous):

My guess is you calculated something incorrectly.

OpenStudy (anonymous):

I am going back over them

OpenStudy (anonymous):

\[g _{7}=\left[\begin{matrix}1 & 0 & 0 & 0 \\ 0 & 1 & 0 & 0.5 \\ 0 & 0 & 1 & 0 \\ 0 & 0 & 0 & 1\end{matrix}\right]\] which would make \[g _{7}^{-1}=\left[\begin{matrix}1 & 0 & 0 & 0 \\ 0 & 1 & 0 & -0.5 \\ 0 & 0 & 1 & 0 \\ 0 & 0 & 0 & 1\end{matrix}\right]\] is that correct?

OpenStudy (anonymous):

Yes I believe so.

OpenStudy (anonymous):

In the notes \[r=\sqrt{x _{\omega}^{2}+y _{\omega}^{2}}\] Should I change r to \[r=\sqrt{x _{\omega}^{2}+y _{\omega}^{2}+z _{\omega}^{2}}\]

OpenStudy (anonymous):

Maybe if it makes sense to you. It depends what it's being used for.

OpenStudy (anonymous):

When I find \[\omega \] I use the quadratic formula, which gives me two values for \[\omega \] I do not know which one to use.

OpenStudy (anonymous):

well both are possible solutions

OpenStudy (anonymous):

In the notes it says just pick one. I took it to mean the as long as \[\omega \] is not complex then both will give you a correct answer. This means that I will only need to do the rest of the calculations with the one I pick and not do it again because the one I picked did not give me the right answer.

OpenStudy (anonymous):

Yes I think you're right. You can just pick one. How are you getting alpha1?

OpenStudy (anonymous):

when I have \[\omega \] and I go to find \[g _{\omega}=g _{1}(\alpha _{1})g _{2}(\alpha _{2})g _{3}(\alpha _{3})g _{4}\] say \[\alpha _{1}\] rotates about the z axis this would mean \[g _{1}(\alpha _{1})=\left[\begin{matrix}1 & 0 & 0 & 0 \\ 0 & \cos (\alpha _{1}) & -\sin (\alpha _{1}) & 0 \\ 0 & \sin (\alpha _{1}) & \cos (\alpha _{1}) & 0 \\ 0 & 0 & 0 & 1\end{matrix}\right]\] and if \[\alpha _{2}\] rotated about the x axis then \[g _{2}(\alpha _{2})=\left[\begin{matrix}\cos (\alpha _{2}) & -\sin (\alpha _{2}) & 0 & 0 \\ \sin (\alpha _{2}) & \cos (\alpha _{2}) & 0 & 0 \\ 0 & 0 & 1 & 0 \\ 0 & 0 & 0 & 1\end{matrix}\right] \] is this the right idea because Ithe manipulator is in x y and z directions

OpenStudy (anonymous):

Yes it's the right idea but you seem to be leaving out the displacements

OpenStudy (anonymous):

\[d _{1}=\left(\begin{matrix}0 \\ 0 \\ 0\end{matrix}\right)\] \[d _{2}=\left(\begin{matrix}0 \\ 0 \\ l _{0}\end{matrix}\right)\] \[d _{3}=\left(\begin{matrix}0 \\ l _{1} \\ 0\end{matrix}\right)\] \[d _{4}=\left(\begin{matrix}0 \\ l _{2} \\ 0\end{matrix}\right)\] but \[g _{4} = \left[\begin{matrix}1 & 0 & 0 & 0 \\ 0 & 1 & 0 & 1 \\ 0 & 0 & 1 & 0 \\ 0 & 0 & 0 & 1\end{matrix}\right]\] because it is just a translation

OpenStudy (anonymous):

why do you say that? g4 is a rotation about the x by alpha4

OpenStudy (anonymous):

Do you say that because you only take the dispalcement part of g4 as part of gw?

OpenStudy (anonymous):

when I find \[g _{\omega} \] using Pieper's approach I break \[g _{4}(\alpha _{4})\] into a translational part \[g _{4}\] and a rotational part

OpenStudy (anonymous):

Yes that's fine. That doesn't mean g_4 is only a translation. It just means you only use the translation part for gw.

OpenStudy (anonymous):

I have class I have to go

OpenStudy (anonymous):

Alright I'll check back tonight.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

back at 6:30

OpenStudy (anonymous):

don't think i'll be on then but i'll check your questions later. Just go over your formulas again. Make sure everything makes sense to you

OpenStudy (anonymous):

I am on #2, I have the polynomials and I need to plot the manipulator movement. Does that mean plot the alpha functions?

OpenStudy (anonymous):

No you should plot the actual manipulator. There was a plot function distributed with homework 2 that can achieve this for you.

OpenStudy (anonymous):

Are you refering to the parametric plot?

OpenStudy (anonymous):

No there is something like planarR3plot or plotPlanarR3

OpenStudy (anonymous):

I found it. It states that I need to plot 5-7 points along the trajectory.

OpenStudy (anonymous):

But it is just a display. Would I show many plots on the same display?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks later

OpenStudy (anonymous):

Thanks. I am on the last question. The demo. We have the spline code made. No errors in running the code. The manipulator starts in the sleep position and moves the home position. We press enter and the screen shows almost a perfect unit matrix with a 17 in the z position (3 4 position). but the manipulator does not move until it finishes and moves back to the sleep position. We have check the code now for about an hour and still can not find anything. Is there a special something we should be lokking for?

OpenStudy (anonymous):

How are you telling the manipulator to follow the trajectory?

OpenStudy (anonymous):

I'm not going to debug your code for you guys. That's your job. My usual technique for debugging is to look at different values throughout the program. Make sure your alpha_traj looks like you would expect it to look etc. The answer to your question is no there is nothing special that you need to look for. There is just a programming error somewhere.

OpenStudy (anonymous):

When this does work, it will move from the home position to the first position. Then only slightly to the next positions until the final position is reached. Then it will move to the sleep position. Is that correct?

OpenStudy (anonymous):

So all we need is the forward kinematics.

OpenStudy (anonymous):

I'm not so sure what forward kinematics has to do with this. You need a video of your manipulator going through a spline trajectory. It should start at the angles specified by alpha_i and end at the angles specified by alpha_f. It will start out moving slowly, speed up in the middle, and slow down toward the end if it is following the proper spline trajectory.

OpenStudy (anonymous):

We had six unknowns and we needed six equations to solve this problem like was done in class. We took the second derivative to come up with the other two equations. Was that correct?

OpenStudy (anonymous):

We used the same method as in the first #2 of this homework to solve for the alphas. how does the manipulator know to speed up and slow down? Is it built into the equations?

OpenStudy (anonymous):

You seem to be confused about what you're doing. You are designing a trajectory (position vs. time) for each alpha. Each alpha has its own cubic trajectory that it should be following. The coefficients of the cubic equation depend on alphai, alphaf, and tf. The purpose of having them follow cubic trajectories is that they slowly speed up at the beginning and slow down at the end. So yes it is built into the equation that the speed is low at the beginning and end but fast in the middle.